1) If (n2+1)t=u + v and (n+1/n)t= u-v, prove that u/v = n+1/n-1 and find t in terms of u,v only

2) For what values of c is x=-3 a root of the equation, x2+(c-1)x-2=c2? What are then the other values of x?


1
Expert's answer
2021-10-21T15:02:14-0400

Solution. 1) Prove that u/v = n+1/n-1.

According to the condition of the problem (n2+1)t=u + v and (n+1/n)t= u-v. Let's add u + v and u-v get


"(n^2+1)t+(n+\frac{1}{n})t=u+v+u-v"

"(n^2+1)t+(\frac{n^2+1}{n})t=2u"

"u=\frac{(n^2+1)(n+1)t}{2n}"

Subtract u-v from u + v get

"u + v -(u -v)=(n^2+1)t-(n+\frac{1}{n})t"

"2v=(n^2+1)t-(\frac{n^2+1}{n})t"

"v=\frac{(n^2+1)(n-1)t}{2n}"

Find the ratio u/v


"\frac{u}{v}=\frac{\frac{(n^2+1)(n+1)t}{2n}}{\frac{(n^2+1)(n-1)t}{2n}}=\frac{n+1}{n-1}."

Find t in terms of u,v. 


"(n+\frac{1}{n})t=u-v \to \frac{(n^2+1)t}{n}=u-v"

On the other hand (n2+1)t=u + v. As result get


"\frac {u+v}{n} = u - v \to n =\frac {u+v}{u-v}"

Substituting n=(u+v)/(u-v) get


"(\frac{(u+v)^2}{(u-v)^2}+1)t=u+v"

"\frac{u^2+2uv+v^2+u^2-2uv+v^2}{(u-v)^2}t=u+v \to \frac{2u^2+2v^2}{(u-v)^2}t=u+v"

"t=\frac{(u+v)(u-v)^2}{2(u^2+v^2)}"

2) If x=-3 a root of the equation, x2+(c-1)x-2=c2 get


"(-3)^2+(c-1)(-3)-2=c^2"

"c^2+3c-10=0"

The roots of the quadratic equation


"\u0441_1=2"

and

"c_2=-5."

For c=2 get equation


"x^2+(2-1)x-2=2^2 \to x^2+x-6=0."

The roots of the quadratic equation "x_1=-3" and "x_2=2" .

For c=-5 get equation


"x^2+(-5-1)x-2=(-5)^2 \to x^2-6x-27=0."

The roots of the quadratic equation "x_3=-3" and "x_4=9".

The other values of x=2 and x =9

Answer. 1) "t=\frac{(u+v)(u-v)^2}{2(u^2+v^2)}" ; 2) "\u0441_1=2" and "c_2=-5"; the other values of x=2 and x =9.